OpenStudy (anonymous):

Solve the equation. Show your work. Check your answer. 3x – 7 = 5x + 19

OpenStudy (haleyelizabeth2017):

Okay so do you know where to start? *hint hint* Try to get the variables on one side and the other numbers on the other.

OpenStudy (haleyelizabeth2017):

There are none to combine

OpenStudy (anonymous):

ok wait nvm

OpenStudy (haleyelizabeth2017):

3x-7=5x+19 -3x -3x combining like terms means they are on the same side of the equal sign.

OpenStudy (haleyelizabeth2017):

lol

OpenStudy (anonymous):

ok but I still got the correct answer witch is -`13

OpenStudy (anonymous):

just made a small mistake

OpenStudy (anonymous):

:)

OpenStudy (haleyelizabeth2017):

2x=-26 x=-13 yeppers lol

OpenStudy (haleyelizabeth2017):

we all make mistakes :)

OpenStudy (anonymous):

Would you know how to check the work?

OpenStudy (haleyelizabeth2017):

plug in the x value in the x spots

OpenStudy (haleyelizabeth2017):

3x-7=5x+19 3(-13)-7=5(-13)+19 -39-7=-65+19 -46=-46 so it is correct